a, b, and c are three positive numbers. If a is 3/5 of c and b is 1/2 of c, then a is what percent greater than b?
- This is from a GRE practice test that does not provide the solution, only the answer. They provided 20 as the answer; however, I can only come up with an answer of 10

Answer by mananth(16949)

a, b, and c are three positive numbers. If a is 3/5 of c and b is 1/2 of c, then a is what percent greater than b?
a= 3c/5
b= c/2
c=2b
c=5a/3
a/b = 6/5
hence 20%